# A Micro-Macro Decomposition-Based Asymptotic-Preserving Random Feature Method for Multiscale Radiative Transfer Equations

### v1

@@ 4 Numerical results
In this section, we present the numerical results of the APRFM for the radiative transfer equation. To enhance the performance of APRFM, in this paper, we introduce a new 1D PoU function $\varphi$ as follows:
$\varphi(z)=\exp(-z^{2}),$ (67)
it can be observed that $\varphi$ is a $C^{\infty}$ function with no compact support. This PoU function is employed in the following numerical examples, except for Example 1. For all subsequent examples, unless explicitly mentioned otherwise, the default activation function is set as the $\tanh$ function. The collocation points are selected as uniform grids, denoted by $(N_{x},N_{v})$ for 1D cases and $(N_{x_{1}},N_{x_{2}},N_{v})$ for 2D cases. The domain partitioning is represented as $(M_{x},M_{v})$ for 1D cases and $(M_{x_{1}},M_{x_{2}},M_{v})$ for 2D cases. All reference solutions in the examples are computed using the finite difference method.
@@ 4.1 One-dimensional problems
with the exact solution $f_{\text{ex}}=1-x$. For the kinetic regime ($\varepsilon=1$), we set $J_{i}^{\rho}=J_{i}^{g}=2^{n-1},n=4,5,6,7$ and $B=2.5$. Other settings are the same as Figure 2 in the previous section. The relative $\ell^{2}$ error of the APRFM solution to the degrees of freedom $Q=2^{n}$ is shown in Figure 5.
Figure 5: Relative $\ell^{2}$ error of the APRFM solution with respect to the degrees of freedom $Q$.
To investigate the dependency of the scale parameter in APRFM on degrees of freedom and collocation points, we consider $\varepsilon=10^{-2^{i}}$, $i=1,2,3,4$ with different random features $J_{i}^{\rho}=J_{i}^{g}=J$ and the collocation points $(N_{x},N_{v})$. We record the relative $\ell^{2}$ error of the APRFM solution with respect to the degrees of freedom and the number of collocation points in Table 3 and Table 4, respectively. We can observe that our APRFM is robust with respect to the scale parameter $\varepsilon$ and needs fewer degrees of freedom and collocation points to achieve high accuracy compared with the RFM in Table 1 and Table 2.
@@
We plot the solution obtained by our APRFM and reference solution in Figure 6. We set $J_{i}^{\rho}=64,J_{i}^{g}=128$ and $B=1.5$. The number of interior collocation points is $(N_{x},N_{v})=(256,128)$, while the number of boundary collocation points is $N_{v}=64$. Additionally, we use $24$ points in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x},M_{v})=(2,4)$, i.e., $M^{\rho}=2,M^{g}=8$. The relative $\ell^{2}$ error of our APRFM is $7.87\times 10^{-3}$ for $\varepsilon=1$ and $5.70\times 10^{-3}$ for $\varepsilon=5\times 10^{-1}$, respectively. It can be observed that the errors of the APRFM are predominantly concentrated at $x=0,1$ and $v=0$.
@@
Figure 6: Plots of $f(x,v)$ in $(x,v)\in[0,1]\times[-1,1]$ by APRFM and finite difference method.
In the kinetic case ($\varepsilon=1$), we also compared our proposed APRFM with the APNN method [28, 27]. As shown in Figure 7, our approach not only achieves slightly lower relative $\ell^{2}$ error compared to the APNN method, but more importantly, it significantly outperforms in terms of computational efficiency, being an order of magnitude faster and requiring considerably fewer parameters. Owing to the uniform accuracy of our APRFM, the comparative results hold similarly for cases where $\varepsilon\ll 1$. This comparative experiment was conducted on a machine equipped with an NVIDIA A800 80G GPU and a 48-core Intel Xeon Gold 6342 CPU. The neural networks for $\rho$ and $g$ in the APNN method are 4-layer fully connected architectures, each with 128 units per hidden layer, utilizing the $tanh$ activation function. We trained the APNN method for 50000 iterations using the Adam optimizer, with a batch size of 8196 for interior points and 2048 for boundary points during the training process.
Figure 7: The comparison of APRFM and APNN.
@@
This case can verify the adaptability of our APRFM under mixed multi-scale conditions. Here, $\varepsilon(x)$ is a function that depends on the spatial variable and smoothly transitions from $\mathcal{O}(10^{-3})$ to $\mathcal{O}(1)$, defined as follows:
@@
We plot the solution obtained by our APRFM and reference solution in Figure 9. We set $J_{i}^{\rho}=96,J_{i}^{g}=128$ and $B=2.5$. The number of interior collocation points is $(N_{x},N_{v})=(512,256)$, while the number of boundary collocation points is $N_{v}=64$. Additionally, we use $32$ points in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x},M_{v})=(2,4)$, i.e., $M^{\rho}=2,M^{g}=8$. The relative $\ell^{2}$ error of our APRFM is $2.88\times 10^{-2}$.
@@ 4.2 Two-dimensional problems
Here, the source function is $G(\bm{x},\bm{v})=(-\cos(\alpha)-\sin(\alpha))\exp(-x-y)/\varepsilon$. In this example, the exact solution is given by $f_{\text{ex}}=\exp(-x-y)$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 10. We set $J_{i}^{\rho}=128,J_{i}^{g}=96$ and $B=3$, with the activation function defined as $\sin(\pi x)$. The number of interior collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,64)$, while the number of boundary collocation points is $(32,64)$. Furthermore, $16$ points are employed in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(2,2,1)$, resulting in $M^{\rho}=4,M^{g}=4$. The relative $\ell^{2}$ error of density function $\rho(\bm{x})$ obtained by our APRFM is $4.92\times 10^{-3}(\varepsilon=1)$ and $4.09\times 10^{-3}(\varepsilon=10^{-3})$.
@@
Here, the source function $G(\bm{x},\bm{v})=1/2$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 11. We set $J_{i}^{\rho}=64,J_{i}^{g}=128$ and $B=0.5$ for $\varepsilon=1$ and $B=1$ for $\varepsilon=10^{-1}$. The number of interior collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,32)$, while the number of boundary collocation points is $(32,32)$. Furthermore, $24$ points are employed in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(1,1,4)$, resulting in $M^{\rho}=1,M^{g}=4$. The relative $\ell^{2}$ error of our APRFM is $3.01\times 10^{-2}(\varepsilon=1)$ and $4.19\times 10^{-2}(\varepsilon=10^{-1})$.
@@
Here, the source function $G(\bm{x},\bm{v})=(-\cos(\alpha)-\sin(\alpha))\exp(-x-y)/\varepsilon$. As before, the exact solution is given by $f_{\text{ex}}=\exp(-x-y)$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 12. We set $J_{i}^{\rho}=128,J_{i}^{g}=96$ and $B=3$ for $\varepsilon=1$ and $B=2$ for $\varepsilon=5\times 10^{-3}$, with the activation function defined as $\sin(\pi x)$. The number of interior collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,64)$, while the number of boundary collocation points is $(32,64)$. Furthermore, $24$ points are employed in the Gauss-Legendre quadrature. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(1,1,4)$, resulting in $M^{\rho}=1,M^{g}=4$. The relative $\ell^{2}$ error of our APRFM is $3.18\times 10^{-4}(\varepsilon=1)$ and $8.74\times 10^{-5}(\varepsilon=5\times 10^{-3})$.
@@
Figure 12: Plots of $\rho(x,y)$ in $(x_{1},x_{2})\in[-1,1]\times[-1,1]$ by APRFM and finite difference method.

### v2

@@ 4 Numerical results
In this section, we present the numerical results of the APRFM for the radiative transfer equation. To validate the effectiveness of our proposed method, we conduct all numerical experiments under the default settings of an initial weight parameter range $B=1$, the number of Gauss-Legendre quadrature points $16$, the partition of unity function
$\psi_{i}(\bm{y})=\prod_{j=1}^{d}\varphi_{b}(\tilde{y}_{ij}),$ (71)
and the activation function $\text{tanh}(\cdot)$. To enhance accuracy, we employ float64 floating-point precision and normalize all PoU functions as follows:
$\tilde{\psi_{i}}(\bm{y})=\psi_{i}(\bm{y})/\sum_{i=1}^{M}\psi_{i}(\bm{y}),$ (72)
ensuring that
$\sum_{i=1}^{M}\tilde{\psi_{i}}(\bm{y})=1.$ (73)
For the sake of simplicity and clarity, we omit all the tildes $\tilde{}$ from $\psi_{i}$ throughout the paper. Second, we upgraded the floating-point precision from float32 to float64. The collocation points are selected as uniform grids, denoted by $(N_{x},N_{v})$ for 1D cases and $(N_{x_{1}},N_{x_{2}},N_{v})$ for 2D cases. The domain partitioning is represented as $(M_{x},M_{v})$ for 1D cases and $(M_{x_{1}},M_{x_{2}},M_{v})$ for 2D cases. In the numerical examples where analytical solutions are not available, the reference solutions are obtained using the finite difference method. We compute the relative $\ell^{2}$ error of the solution $f(x,v)$ over all uniform grids $\{(\bm{x}^{i},\bm{v}^{i})\}_{1\leq i\leq I}$ with mesh size $(N_{x},N_{v})=(128,256)$ for 1D cases and $(N_{x_{1}},N_{x_{2}},N_{v})=(64,64,32)$ for 2D cases:
$\text{error}:=\sqrt{\frac{\sum_{i=1}^{I}|f^{\text{approx}}(\bm{x}^{i},\bm{v}^{i})-f^{\text{ref}}(\bm{x}^{i},\bm{v}^{i})|^{2}}{\sum_{i=1}^{I}|f^{\text{ref}}(\bm{x}^{i},\bm{v}^{i})|^{2}}},$ (74)
where $f^{\text{approx}}$ denotes the solution approximated by vanilla RFM or APRFM, and $f^{\text{ref}}$ represents the reference solution.
@@ 4.1 One-dimensional problems
with the exact solution $f_{\text{ex}}=1-x$. For the kinetic regime ($\varepsilon=1$), we set $J^{\rho}=J^{g}=2^{n-1},n=3,4,5,6,7$ and $(M_{x},M_{v})=(1,1)$. Other settings are the same as Figure 2 in the previous section. The relative $\ell^{2}$ error of the APRFM solution to the degrees of freedom ${Z}=2^{n}$ is shown in Figure 5.
Figure 5: Relative $\ell^{2}$ error of the APRFM solution with respect to the degrees of freedom.
To investigate the dependency of the scale parameter in APRFM on degrees of freedom and collocation points, we consider $\varepsilon=10^{-2^{i}}$, $i=1,2,3,4$ with different random features $J^{\rho}=J^{g}=J$ and the collocation points $(N_{x},N_{v})$. We record the relative $\ell^{2}$ error of the APRFM solution with respect to the degrees of freedom and the number of collocation points in Table 4 and Table 5, respectively. We can observe that our APRFM is robust with respect to the scale parameter $\varepsilon$ and needs fewer degrees of freedom and collocation points to achieve high accuracy compared with the RFM in Table 1 and Table 2.
@@
We plot the solution obtained by our APRFM and reference solution in Figure 6. We set $J^{\rho}=64,J^{g}=128$. The number of interior collocation points is $(N_{x},N_{v})=(128,256)$. Additionally, the domain is partitioned as $(M_{x},M_{v})=(2,4)$, i.e., $M^{\rho}=2,M^{g}=8$. The relative $\ell^{2}$ error of our APRFM is $7.08\times 10^{-3}$ for $\varepsilon=1$ and $5.56\times 10^{-3}$ for $\varepsilon=5\times 10^{-1}$, respectively. It can be observed that the errors of the APRFM are predominantly concentrated at $x=0,1$ and $v=0$.
@@
Figure 6: Plots of $f(x,v)$ in $(x,v)\in[0,1]\times[-1,1]$.
In the kinetic case ($\varepsilon=1$), we also compared our proposed APRFM with the finite difference method (FDM) and APNN method [28, 27]. As shown in Figure 7, our approach not only achieves lower relative $\ell^{2}$ error compared to the APNN method, but more importantly, it significantly outperforms in terms of computational efficiency, being an order of magnitude faster and requiring considerably fewer parameters. Owing to the uniform accuracy of our APRFM, the comparative results hold similarly for cases where $\varepsilon\ll 1$. For this 1D example, our method does not provide a significant advantage in terms of computation time compared to the traditional finite difference method. In this example, due to the large number of collocation points and the extensive partitioning of the domain, the assembly of the matrix took 144.4 seconds, accounting for $88\%$ of the total computation time. However, it is important to highlight that our APRFM demonstrates notable improvements in computation time for 2D cases, as detailed in Example 4. This comparative experiment was conducted on a machine equipped with an NVIDIA A800 80G GPU and a 48-core Intel Xeon Gold 6342 CPU. The neural networks for $\rho$ and $g$ in the APNN method are 4-layer fully connected architectures, each with 128 units per hidden layer, utilizing the $\tanh$ activation function. We trained the APNN method for 50000 iterations using the Adam optimizer, with a batch size of 8196 for interior points and 2048 for boundary points during the training process. The number of inner-layer network parameters in our APRFM that do not participate in training is 1152. Other settings are the same as those in Example 2.
Figure 7: The comparison of APRFM and APNN in 1D case.
@@
This case can verify the adaptability of our APRFM under mixed multi-scale conditions. Here, $\varepsilon(x)$ is a function that depends on the spatial variable and smoothly transitions from $\mathcal{O}(10^{-2})$ to $\mathcal{O}(1)$, defined as follows:
@@
For this case, we briefly derive the corresponding micro-macro decomposition system. First, we decompose $f(x,v)$ as
$f(x,v)=\rho(x)+\varepsilon(x)g(x,v),$ (79)
with $\left\langle g\right\rangle=0$. Substituting this into equation (77), we rewrite it as
$v\cdot\nabla_{x}(\rho(x)+\varepsilon(x)g(x,v))+g(x,v)=0.$ (80)
Applying the orthogonal projection operators $\Pi:\Pi(\cdot)(v)=\left\langle\cdot\right\rangle$ and $\text{Id}-\Pi$ to equation (80), we derive the micro-macro system for this mixed-scale problem:
$\begin{cases}\left\langle v\cdot\nabla_{x}(\varepsilon(x)g)\right\rangle=0,\\ v\cdot\nabla_{x}\rho+(\text{Id}-\Pi)(v\cdot\nabla_{x}(\varepsilon(x)g))+g=0.\end{cases}$ (81)
Thus, we can employ our APRFM to solve the micro-macro system (81).
We plot the solution obtained by our APRFM and reference solution in Figure 9. We set $J^{\rho}=64,J^{g}=128$. The number of collocation points is $(N_{x},N_{v})=(128,256)$. The domain is partitioned as $(M_{x},M_{v})=(2,4)$, i.e., $M^{\rho}=2,M^{g}=8$. The relative $\ell^{2}$ error of our APRFM is $1.72\times 10^{-2}$.
@@ 4.2 Two-dimensional problems
Here, the source function $G(\bm{x},\bm{v})=(-\cos(\alpha)-\sin(\alpha))\exp(-x_{1}-x_{2})/\varepsilon$. In this example, the exact solution is given by $f_{\text{ex}}=\exp(-x_{1}-x_{2})$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 10. We set $J^{\rho}=32,J^{g}=32$ and the number of collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,64)$. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(1,1,1)$, i.e., $M^{\rho}=1,M^{g}=1$. The relative $\ell^{2}$ error of density function $\rho(\bm{x})$ obtained by our APRFM is $3.48\times 10^{-4}(\varepsilon=1)$ and $2.28\times 10^{-4}(\varepsilon=10^{-3})$. For $\varepsilon=1$ of this example, the computation time for our APRFM was 6.29 seconds, while the computation time for the finite difference method was 111.20 seconds. It should be noted that as the number of collocation points, partition regions, and random feature functions increase, the computation time for the APRFM will correspondingly increase.
@@
Here, the source function $G(\bm{x},\bm{v})=1/2$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 11. We set $J^{\rho}=64,J^{g}=128$ and the number of interior collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,32)$. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(1,1,4)$, resulting in $M^{\rho}=1,M^{g}=4$. The relative $\ell^{2}$ error of our APRFM is $3.42\times 10^{-2}(\varepsilon=1)$ and $4.43\times 10^{-2}(\varepsilon=10^{-1})$. Additionally, we have documented the relative $\ell^{2}$ errors corresponding to different PoU configurations in Table 6. In this case, it has been observed that simply increasing the number of partitions does not necessarily lead to better results.
@@
Table 6: The dependency of the RFM on PoU. The collocation points are uniform grids with $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,32)$ and the number of random feature functions $J^{\rho}=64,J^{g}=128$.
$\varepsilon$ $(M_{x_{1}},\,M_{x_{2}},\,M_{v})$
(1, 1, 1)111(1,\,1,\,1)( 1 , 1 , 1 ) (1, 1, 2)112(1,\,1,\,2)( 1 , 1 , 2 ) (1, 1, 4)114(1,\,1,\,4)( 1 , 1 , 4 ) (1, 1, 8)118(1,\,1,\,8)( 1 , 1 , 8 )
[…]
@@
Here, the source function $G(\bm{x},\bm{v})=(-\cos(\alpha)-\sin(\alpha))\exp(-x_{1}-x_{2})/\varepsilon$. As before, the exact solution is given by $f_{\text{ex}}=\exp(-x_{1}-x_{2})$. We plot the density function $\rho(\bm{x})$ obtained by our APRFM and reference solution in Figure 12. We set $J^{\rho}=64,J^{g}=128$ and the number of collocation points is $(N_{x_{1}},N_{x_{2}},N_{v})=(32,32,64)$. The domain is partitioned as $(M_{x_{1}},M_{x_{2}},M_{v})=(1,1,4)$, resulting in $M^{\rho}=1,M^{g}=4$. The relative $\ell^{2}$ error of our APRFM is $7.54\times 10^{-7}(\varepsilon=1)$ and $1.60\times 10^{-6}(\varepsilon=5\times 10^{-3})$. Besides, we replaced the activation function with $\sin(\pi x)$, while keeping all other settings unchanged. The corresponding density function $\rho$ is shown in Figure 13, with the relative $\ell^{2}$ error of $7.34\times 10^{-7}(\varepsilon=1)$ and $4.50\times 10^{-6}(\varepsilon=5\times 10^{-3})$.
@@
Figure 12: Plots of $\rho(x,y)$ in $(x_{1},x_{2})\in[-1,1]\times[-1,1]$.
(a) Reference solution v.s. APRFM solution ($\varepsilon=1$).
Figure 13: Plots of $\rho(x,y)$ in $(x_{1},x_{2})\in[-1,1]\times[-1,1]$.